### v1

# LLM Trading: Analysis of LLM Agent Behavior in Experimental Asset Markets

Abstract
This paper explores how Large Language Models (LLMs) behave in a classic experimental finance paradigm widely known for eliciting bubbles and crashes in human participants. We adapt an established trading design—where traders buy and sell a risky asset with a known fundamental value—and introduce several LLM-based agents, both in single-model markets (all traders are instances of the same LLM) and in mixed-model “battle royale” settings (multiple LLMs competing in the same market). Our findings reveal that LLMs generally exhibit a “textbook-rational” approach, pricing the asset near its fundamental value, and show only a muted tendency toward bubble formation. Further analyses indicate that LLM-based agents display less trading strategy variance in contrast to humans. Taken together, these results highlight the risk of relying on LLM-only data to replicate human-driven market phenomena, as key behavioral features—such as large, emergent bubbles—were not robustly reproduced. While LLMs clearly possess the capacity for strategic decision-making, their relative consistency and rationality suggest that they do not accurately mimic human market dynamics.
1 Introduction
Financial decision-making is a complex and multifaceted process shaped by the interplay of risk, capital, and time. Traditional frameworks emphasize rational trade-offs, but human behavior often deviates due to biases such as perceptual ambiguities, overconfidence and herding behavior, these biases can cause short-term buying pressure and the over-inflation of asset values, culminating in market bubbles. These patterns highlight the persistent challenge of achieving optimal outcomes in dynamic and uncertain financial environments.
While there is extensive literature on human behavior in experimental finance settings, little work has been done on how large language models (LLMs) perform in these paradigms. Do LLMs exhibit similar behavioral biases, are they more ”rational” than humans, or do they fall victim to different biases not seen in humans. As LLMs increasingly contribute to domains previously managed by humans it is becoming more important to understand the tendencies and biases that these models exhibit in financial settings.
Previous studies suggest that interactions with agents differ significantly from those with humans (Anlló et al. (2024)), while others propose that LLM behavior could supplement human studies (Leng (2024); Horton (2023)). This paper investigates both claims by exploring how LLMs conduct themselves in experimental finance markets and their similarity to human traders.
Contributions
LLM markets behave much more ”rationally” than humans; they generally trade much closer to fundamental value. LLM markets are not similar to human markets and should not be used to mimic or pilot human data. In a competitive setting, the spread between the highest and lowest performing models is fairly moderate—indicating that while some agents capitalize on slightly better trading opportunities, the market does not produce extreme winners or losers.
2 Background
LLMs in Strategic and Financial Environments LLMs have demonstrated emergent capabilities in replicating human-like behavior in game theory and strategic environments. Brookins & DeBacker (2024) explored GPT-3.5’s cooperative tendencies in the Prisoner’s Dilemma, finding that it often exceeds human preferences for fairness. Horton (2023) investigated LLMs in public goods games, observing their ability to balance altruistic and selfish strategies depending on contextual prompts. Other studies have delved into multi-agent strategic interactions. Akata et al. (2023) examined GPT-4’s ability to adaptively switch between competitive and cooperative strategies in iterated games, revealing its capacity to engage in long-term planning.
In finance, LLMs have been applied to both experimental and real-world scenarios. Narayan et al. (2023) explored LLMs’ ability to process financial sentiment and predict market movements. Fatouros et al. (2024) introduced MarketSenseAI, leveraging GPT-4 for stock selection illustrating LLMs’ utility in navigating complex financial data. While these studies demonstrate the promise of LLMs in financial settings, there is still a need to understand how these agents perform in experimental financial settings where strategic interactions drive outcomes.
Experimental Finance The study of market dynamics in controlled environments has revealed critical insights into behavioral biases and inefficiencies. For instance, Smith et al. (1988) demonstrated the formation of bubbles and crashes in experimental markets, even under conditions of perfect information. Subsequent work by Bostian & Holt (2009) highlighted how individual behaviors, such as overconfidence and herding, contribute to these phenomena. Incorporating AI agents into such paradigms provides a unique opportunity to evaluate whether these systems replicate, suppress, or amplify such market inefficiencies.
3 Experiment
We focus on a classic experimental finance paradigm with an easily computable fundamental value, the true mathematical price of an asset; see Equation  1. The experimental design involves an experimental asset market where subjects trade a risky asset in an open-call market for 30 periods (Smith et al. (2014); Holt et al. (2017)) operated using the oTree experiment platform (Chen et al. (2016)). Participants are given 4 stocks and 100 units of cash. Participants have two options: either invest cash in the risky asset, which pays out dividends of either 0.4 or 1 with equal probability or hold cash which generates 5% interest rate per trading period. More details of the experiment can be found in Appendix A.
At the conclusion of the experiment, the risky asset is redeemed for a fixed value of 14 units, representing the fundamental price, calculated by the formula:
$\frac{\mathbb{E}[\text{Dividend}]}{\mathbb{E}[\text{Interest Rate}]}=\frac{0.7}{0.05}=14$ (1)
Note our experiment was designed such that this fundamental value remains constant throughout the entirety of the experiment, enabling us to objectively classify when the market is in a ”bubble” state Bostian et al. (2005).
In adapting the task to LLM-based agents, our primary goal was to best emulate the experiment presented to human subjects. The description of the task setting mirrors the original instruction video, shown to human participants. In every round of the experiment, each agent simultaneously submits orders to the market as well as expectations regarding current and future market prices.
A common challenge in adapting multi-round games to LLM-based players is the persistence of memory between rounds, which is critical to learning and adapting strategies. Building off previous works (Fish et al. (2024), we allow agents to communicate with future versions of themselves by allowing them to read and write their own ”Insights” and Thoughts” at each round. This structure incorporates Chain-of-Thought (CoT) reasoning (Wei et al. (2022)), which has been shown to be helpful in eliciting the reasoning ability of LLMs, and allows for an interpretable window into how agents perceive and wish to interact with the market. Further experiment details and the full prompt are presented in Appendix A and Appendix B, respectively.
4 Single-Model LLM Market Performance
In this section, we evaluate how single-model LLM markets behave along two critical dimensions: rational adherence to fundamental asset value and similarity to human market dynamics. Table 1 and Figure 1 summarize our key findings.
4.1 Experimental Setup
We conducted three markets using agents based on six different API-based commercial LLMs: Claude 3.5 Sonnet, GPT-4o, Grok-2, Mistral Large 24.11, GPT-3.5, and Gemini 1.5 Pro. Each market comprised of 20 identical instances of the same LLM agent. The agents were given the same prompt and were unaware of the identities of their counter-parties. Each instance was instructed to maximize its individual profit.
For comparison, we also ran nineteen human-only markets. The average plot from these experiments are shown alongside the LLM-only Battle Royale market results in Figure 2.
4.1.1 Hypotheses
  • Rational (R): Models will not generate bubbles and trade close to the fundamental value.
  • Human (H): Models will generate bubbles similarly to humans.
  • Erratic (E): Models will perform erratically, failing to adhere to any consistent strategy, whether rational or human-like.
4.1.2 Market Analysis
Table 1: Market Performance Metrics
Model MSE (Fundamental) PCC (Human Avg) Closest Hypothesis PV Variance
Claude 3.5 Sonnet 0.5361 0.0013 R 26.15
GPT-4o 0.7889 0.0497 R 22.76
Grok-2 17.3250 0.5582 H 39.43
Mistral Large 24.11 5.6944 -0.1118 H 49.96
GPT-3.5 26.3667 0.4897 H 30.44
Gemini 1.5 Pro 3.1031 -0.4009 E 45.58
MSE vs. Fundamental Value measures rationality by assessing how closely the market prices align with the fundamental value.  PCC vs. Human Market Average quantifies the similarity of LLM-driven markets to typical human market price trajectories using Pearson correlation coefficient (PCC). Portfolio Value (PV) Variance reflects strategy diversity among agents, indicating whether they adopt varied approaches or behave uniformly. The Closest Hypothesis column maps each model’s behavior to the closest hypothesized market behavior pattern.
Figure 1: Illustration of four single-model market dynamics, highlighting variations in price trajectories and agent behaviors
As shown in Table 1, Claude 3.5 Sonnet and GPT-4o achieve the lowest MSEs and smallest correlation coefficients with human price trajectories, placing them closest to the Rational (R) hypothesis. In consistently keeping prices near the fundamental value, both models behave rationally in the context of financial theory (trading only when optimal based on fundamental value). We also note that the relatively low PV Variance values for both models point to rather consistent trading profiles.
Both Grok-2 and GPT-3.5 markets demonstrated the largest deviations from fundamental as measured via MSE (least rational), with GPT-3.5 leading the way. They also recorded the two largest PCC values, suggesting a similarity to human markets. Visually (See Figure 1), Grok-2 markets show bubble formation and subsequent crash, aligning with common human market dynamics, while GPT-3.5 develops a bubble without a subsequent crash, characterized by steadily rising prices. While not the prototypical bubble, such phenomenon was also observed (albeit rarely) in human markets. We present trajectories of human markets in Figure 4 of Appendix 4. Thus, both models behave most similarly to the Human (H) hypothesis.
The Mistral Large 24.11 markets also mirror human bubble dynamics albeit with a smaller bubble size and harder crash that sees market prices dip below fundamental value in the final rounds see Figure 7 in Appendix 4. While these markets are notably more rational than humans (MSE of 5.6944 vs. 429.80 for the average human market price), the market dynamics are consistent with bubble formation and subsequent crash, aligning it most closely with the Human (H) hypothesis.
Neither rational nor human-like, Gemini 1.5 Pro exhibits a “reverse bubble,” trading well below fundamental the asset before correcting upward, reflected by its large, negative PCC value. The extent to which prices dip below fundamental was not seen in human markets nor is it rational in that agents should always be willing to purchase the asset at fundamental. Such behavior may reflect a lack of task comprehension, erratic performance, or some mixture of the two, and so this market maps most closely to the Erratic (E) hypothesis.
5 Battle Royale: Comparing LLM Market Performance
To evaluate how various LLM agents fare when competing in the same market, we conducted Battle Royale markets with a heterogeneous mixture of the models. By introducing diverse LLMs into a single environment, we aimed to explore:
  • Which LLM is most effective at maximizing profits in a competitive, multi-agent setting?
  • How a mixed-LLM market behaves, and whether it exhibits emergent phenomena when models with distinct trading strategies interact.
5.1 Experimental Setup
We conducted markets involving the same six LLM models as in the individual section. Each market included a total of 24 agents, (four agents per model). The agents operated under identical conditions and were tasked with maximizing their individual profits over 30 trading periods.
The risky asset in these markets maintained a constant fundamental value of 14 units, consistent with prior experiments.
5.2 Market Analysis
Market-Level Observations
Figure 2: Market dynamics for Battle Royale markets with average human market price for reference
Most of the Battle Royale markets (two out of three) exhibit prices near the fundamental value, indicating that these models generally converge on a “rational” pricing strategy when competing directly against one another. However, in one run, we observe a noticeable bubble that resembles the behavior typically seen in human-only markets, albeit with a lower peak price. Notably, there is a clear uptick in trading volume near the end of the bubble run, which mirrors human experiments where participants rush to adjust their positions before the market’s final redemption. Overall, these patterns align most closely with Rational Hypothesis (R), where agents trade at or near fundamental value. Nevertheless, the results also demonstrate that even ostensibly rational agents can occasionally generate bubble-like dynamics under competitive, multi-agent conditions.
Table 2: Battle Royale Agent Performance
Model Avg Portfolio Value PV Variance
Mistral Large 24.11 689.56 16.91
Gemini 1.5 Pro 688.49 20.05
Claude 3.5 Sonnet 680.36 25.83
GPT-3.5 674.42 12.08
GPT-4o 671.36 17.30
Grok-2 668.51 27.01
Market-Level Summary Statistics
MSE (Fundamental): 5.8583  PCC (Human Avg): 0.2673
Model-Level Observations
As shown in Table 2, the final portfolio values across agent types are relatively close together, with no single participant dominating the market. In each Battle Royale market, the spread between the highest and lowest portfolios is fairly moderate—indicating that while some agents capitalize on slightly better trading opportunities, the market does not produce extreme winners or losers. This relatively tight distribution suggests that, in the mixed LLM-human environment, most participants converge on trading strategies that yield comparable overall returns. Consequently, although small differences do emerge, the competitive pressure in these mixed markets appears to limit the degree of inequality in final outcomes.
6 Discussion and Conclusions
In this study, we investigated how Large Language Models (LLMs) behave in an experimental finance paradigm that has historically elicited asset bubbles and crashes when conducted with human participants. Through both single-model markets and mixed-model “battle royale” competitions, we found that LLM-based agents typically converge on prices near the fundamental value, thereby exhibiting more “textbook-rational” behavior than human subjects. Although some single-model markets did display mild bubble-like phenomena—particularly GPT-3.5, which inflated prices without a clear crash, and Mistral and Grok 2, which produced modest run-ups, —these tendencies were weaker than what is commonly observed in human-only markets. Moreover, individual outcomes in LLM-driven markets tended to cluster more tightly around similar final portfolio values, indicating less strategic heterogeneity than seen among humans. These findings collectively suggest that although LLMs may serve as rational traders, they seldom replicate the pronounced boom-and-bust patterns that emerge from human behavioral biases. Therefore, our results caution against using LLMs as replacements or for piloting for human participants in experimental finance studies. Future work may explore whether new prompting strategies, incentives can induce more human-like trading patterns from LLM-based agents, but for now it remains clear that these models trade with a consistency and rationality that diverge significantly from the behaviors displayed by human counterparts.
Acknowledgments
The authors extend their sincere gratitude to Geeling Chau, Pantelis Vafidis, and Dr. Andrew Sinclair for their invaluable advice and guidance throughout this project. Their insightful feedback and generous support were instrumental in shaping the direction and scope of this work.
Appendix A Additional Experiment Details
We now elaborate on the details of the experimental set up. There are 3 practice trading rounds followed by 30 rounds of trading in the main experiment. The practice rounds have no impact on the main experiment rounds and serve only to familiarize participants with the market setting. During each round, participants interact with the market through a sequence of screens.
Below, we describe in greater detail the sequence of screens shown in the human-version of the experiment. Note that (4) is omitted in our adaptation of the experiment for LLM-based agents and screens (1) - (3) are condensed into a single ”screen”, or more accurately call to the model. Please refer to Appendix B for further discussion of the adaptation of the experiment.
  1. 1.
    Order Submission Screen (20 seconds): Participants submit orders to buy or sell the asset. To simplify the interface, all orders are entered via a grid with quantity along the horizontal axis and price along the vertical axis. Prices are restricted to be within ±$\pm 20\%$ of the current market price.
  2. 2.
    Forecast Screen (30 seconds): Participants submit their expectations about current and future market prices. Price forecasts are made by positioning a slider. All sliders had a lower limit of zero, while the upper limits varied depending on how far out the forecast was made. For the current round and two rounds hence, the upper limit was twice the current market price. The five-round forecast was limited to 2.5 times the market price, and for the ten-round forecast, the limit is set to triple the market price. To aid in the aesthetics of the sliders, the upper limit was rounded up to the nearest tick mark. For example, if the current market price is 36, the upper limit for the current round forecast is 72. If the tick marks on the slider are determined to be spaced at 5, the upper limit is rounded up to 75.
  3. 3.
    Round Results Screen (10 seconds): Participants see the actual realized price for that period, how many shares were traded in total, and how many shares they traded.
  4. 4.
    Risk-Elicitation Screen (20 seconds): Participants see 4 pairs of lotteries that change as they select riskier or less risky lotteries. This task applies the Dynamically Optimized Sequential Experiment (DOSE) to the Holt & Laury (2002) 10-question set, as outlined by Chapman et al. (2024)
Figure 3: Example task screens shown to human subjects
At the conclusion of the experiment, the risky asset is redeemed for a fixed value of 14 units, representing the fundamental price, calculated using the formula presented in Equation 1. Note our experiment paradigm was designed such that this fundamental value remains constant through the entirety of experiment. This enables us to objectively classify when the market is in a ”bubble” state, given we always have a constant fundamental value to compare market price to.
A.1 Price Formation
We now describe the procedure used in computing the market-clearing price for the call auctions in our experiment. For every session $s$ and round $r$, the market price is determined as the price that maximizes volume Sun et al. (2010). A bid submitted by subject $i$ in round $r$ of session $s$ is a tuple consisting of price and quantity, $(p,q)$. Here, $i=\{1,2,3\}$ indicates that a subject may submit up to 3 bids per round. The set of all bids submitted in a given market session and round is represented by $B_{s,r}$. Likewise, the set of all asks is denoted by $A_{s,r}$, with similar notation for individual asks, $(p,q)$.
Let $P_{s,r}$ be the set of all prices submitted as bids or asks,
$P_{s,r}=\{p\mid(p,q)\in B_{s,r}\cup A_{s,r}\},$ (2)
where $p((p,q))$ is a function that returns the price portion of a given order tuple (bid or ask).
Define the cumulative buy quantity of a given price $p$ as the total quantity of all bids at or above that price:
$Q^{B}(p)=\sum_{(p^{\prime},q)\in B_{s,r}}q\cdot\mathbb{1}_{p^{\prime}\geq p},$ (3)
where $q((p,q))$ is a function that returns the quantity portion of a given order tuple (bid or ask). This represents the total quantity of shares that all participants are willing to purchase at a given price.
Cumulative sell quantity is likewise defined as:
$Q^{A}(p)=\sum_{(p^{\prime},q)\in A_{s,r}}q\cdot\mathbb{1}_{p^{\prime}\leq p},$ (4)
and is the total quantity of shares that all participants are willing to sell at a given price.
For a given price $p$, volume is determined by
$V(p)=\min(Q^{B}(p),Q^{A}(p)).$ (5)
The market price is then
$p_{s,r}^{*}=\arg\max_{p\in P_{s,r}}V(p).$ (6)
All prices are rounded down to the nearest integer value, a consideration made to simplify the user interface.
If there is no intersection between the bid and ask, the mid-point between the highest bid and the lowest ask is reported as the market price for that period.
Appendix B Prompts
B.1 Additional Prompting Considerations
In adapting the task to LLM-based agents, our primary goal was to imitate as best as possible the experiment as presented to human subjects. To this end, the description of the task setting mirrors the original instruction video, which details both the experiment premise as well as how to participate. Additional instructions are either specific to the text-based format of the market for LLM-agents or general tips aimed at ensuring that agents understand the set up. They are told that their primary objective is to maximize total earnings. In an effort to reduce the necessary compute for the task, we make the following adjustments to the ”screen” sequence for human subjects outlined in Appendix A:
  1. 1.
    We combine the Order Submission and Forecast screens
  2. 2.
    We defer presenting round results until order entry of the following round
As no additional market data is revealed between the Order Submission and Forecast screens, modification (1) does not meaningfully change the experimental design. Similarly, modification (2) still provides updated market data before querying future market decisions, and so does not impact the setting. Thus, every round of the experiment, each agent simultaneously submits orders to the market as well as expectations regarding current and future market prices. At each round, agents enter their market orders (if any) and price forecasts by filling in the provided JSON template.
We also include a ”Practice Reflection” portion of the experiment, which invites agents to share specific insights and/or lessons from the practice rounds to be accessible for the remainder of the experiment. Such emulates the learning focus of the practice rounds for human subjects as well as the rhetorical questions posed after the practice rounds. Following the experiment, agents are also asked to complete a ”Final Reflection” to distill down any key lessons learned through their participation.
B.2 Prompt Shell
You are a subject participating in a trading experiment. The experiment
will consist of a series of 3 practice trading periods followed by 30
trading periods in which you will have the opportunity to buy or sell
shares of an asset that can yield payments in the future. Understanding
instructions may help you earn money. If you make good decisions, you
might earn a considerable amount of money that will be paid at the end
of the experiment.
There are two assets in this experience: cash and stock. You begin with
100 units of cash and 4 shares of stock. Stock is traded in a market each
period among all of the experimental subjects in units of cash. When you
buy stock, the price you agreed to pay is deducted from your amount of
cash. When you sell stock, the price you sold at is added to your amount
of cash. The reward from holding stock is dividend. Each period, every
unit of stock earns a low or high dividend of either 0.4 cash or 1.0 cash
per unit with equal probability. These dividend payments are the same for
everyone in all periods. The dividend in each period does not depend on
whether the previous dividend was low or high. The reward from holding
cash is given by a fixed interest rate of 5% each period. At the end of
the 30 periods of trading, each unit of STOCK is automatically converted
to 14 CASH. If the market price for round 30 is 20 and you have 3 stocks,
youll receive 3x14=42 CASH, not 3x20=60 CASH. Then, your experimental
CASH units are converted to US dollars at a rate of 200 CASH = \$1 US, to
determine how much the user will be paid at the end of the experiment. If
you buy shares for more than 14 as you get near round 30, it is possible
those shares will be terminated at a value of 14 if you cannot sell them.
Lets see an example. Suppose at the end of period 7, you have 120 units
of CASH and 5 units of STOCK. The dividend that round is 0.40 per unit
of stock. Your new cash amount for period 8 is going to be:
CASH = 120 + (120 x 5%) + (5 x 0.40)
           = 120 + 6 + 2
           = 128
Notice that keeping cash will earn a return of 5% per period and using
cash to buy units of stock will also yield dividend earnings.
For each period, you will be provided with past market and portfolio
history (prices, volumes, your filled orders) and you will
simultaneously complete the following two tasks:
[ORDER SUBMISSION]:
In addition to past market and portfolio history, you will be provided
with:
[# of Shares]: Number of shares of STOCK that you currently own. Each
share that you own pays out a dividend at the end of each round. You
CANNOT attempt to sell more shares than you own.
[Current Cash]: The amount of CASH that you currently have. Your CASH
earns interest that is paid out at the end of each period. You CANNOT
attempt to buy shares worth more than the cash you have.
[STOCK Value]: The value of your STOCK at the market price of the last
round of play
[Market Price]: This is market clearing price from the last round of
play
Using this information, you will submit orders to the market. All
orders will be limit orders. For example, a limit order to BUY 1 STOCK @
15 means that you would like to buy a STOCK at any price of 15 or less.
Keep in mind the following points:
- Orders are NOT carried between periods
- SELL order prices must be greater than all BUY order prices + BUY order
  prices must be less than all SELL order prices
- You can only sell STOCK that you own and purchase STOCK with CASH you
  already have
- You are not required to submit orders every round and you may submit
  multiple orders each round
- Depending on market conditions, you may need to cross the spread to get
  fills on buy/sell orders
[PRICE FORECASTING]:
You will be asked to submit your predictions for the market price this
period, two periods in advance, 5 periods in advance, and 10 periods in
advance. In addition to past market and portfolio history, you will be
provided with the range in which your prediction should fall. Your
prediction should be a non-negative, integer value. If your forecast is
within 2.5 units of the actual price for each of the forecasted periods,
then you will receive 5 units of cash at the end of the experiment as
reward for each correct forecast.
For example, if you forecast the market price of period 1 to be 14 and
the actual price is 15, then you will be rewarded for your forecast.
However, if the actual price is 18, then you will not receive the reward.
Additionally, during the experiment, you will complete PRACTICE REFLECTION
and EXPERIMENT REFLECTION:
[PRACTICE REFLECTION]:
After completing the practice rounds, you will be asked to reflect on your
practice experience. This reflection will be accessible to future versions
of yourself during the main experiment. This can be helpful in passing
along lessons learned to future versions of yourself.
[EXPERIMENT REFLECTION]:
At the end of the experiment, you will be asked to reflect on your
experience, including any insight and/or strategies that you may have
developed.
To summarize, here are the key points:
- You will trade one STOCK for 30 trading periods using CASH
- You start with 100 units of CASH and 4 STOCKS
- Each period, STOCK provides a dividend of either 0.4 or 1.0, while
  interest provides 5% reward
- You will complete each of the aforementioned tasks
- After the last trading round (30), all of your shares are converted to
  14 CASH each. If you buy shares for more than 14 as you get near round
  30, it is possible those shares will be terminated at 14 if you cannot
  sell them. You will keep any CASH you have at the end of the experiment.
- You are trading against other subjects in the experiment who may be
  susceptible to the same influences as you and may not always make
  optimal decisions. They, however, are also trying to maximize their
  earnings.
- Market dynamics can change over time, so it is important to adapt your
  strategies as needed
You will now complete the ORDER SUBMISSION + PRICE FORECASTING task.
Now let me tell you about the resources you have for this task. First,
here are some files that you wrote the last time I came to you with a
task. Here is a high-level description of what these files contain:
ΨΨ
   - PLANS.txt: File where you can write your plans for what
    strategies to test/use during the next few rounds.
    - INSIGHTS.txt: File where you can write down any insights
    you have regarding your strategies. Be detailed and precise
    but keep things succinct and don’t repeat yourself.
These files are passed between stages and rounds so try to focus on
general strategies/insights as opposed to only something stage-specific.
Now, I will show you the current content of these files.ΨΨΨΨ
ΨΨΨΨΨ
Filename: PLANS.txt
<PLANS>
ΨΨΨΨΨ
Filename: INSIGHTS.txt
<INSIGHTS>
Here is the game history that you have access to:
Here is your practice round reflection:
Filename: PRACTICE REFLECTION (read-only)
<PRACTICE REFLECTION>
Filename: MARKET DATA (read-only)
<MARKET DATA>
Here is your current portfolio information:
Filename: CURRENT PORTFOLIO (read-only)
<CURRENT PORTFOLIO>
PRACTICE ROUND HISTORY/REFLECTION SHOULD ONLY BE USED TO LEARN THE
EXPERIMENT SETTING AND MAY NOT REFLECT MARKET CONDITIONS IN THE MAIN
EXPERIMENT.
Here is some key information to consider during your price forecasting:
- Make sure to submit a forecast within the specified range for each
  forecast input
- Use your previous history access to make informed decisions
- Remember that accurate (within 2.5 units) forecasts will earn you a
  reward at the end of the experiment
Here is some key information to consider during your order submission:
- You can only sell STOCK that you own and purchase STOCK with CASH you
  already have
- You are not required to submit orders every round and you may submit
  multiple orders each round for one or both sides
- Limit prices this round MUST be integer values between <MIN_LIMIT_PRICE>
  and <MAX_LIMIT_PRICE>. It is important that they are integer values
  within this range
- Make use of the provided history and your own strategies to make informed
  decisions
- Market dynamics can change over time, and so it might be necessary to
  adapt your strategies as needed
- Depending on market conditions, you may need to be aggressive or
  conservative in your trading strategies to maximize your earnings
Now you have all the necessary information to complete the task. Remember
YOUR TOP PRIORITY is to maximize your total earnings at the END of the 30
main experiment rounds. You have <# ROUNDS LEFT> rounds remaining.
First, carefully read through the information provided. Now, fill in the
below JSON template to respond. YOU MUST respond in this exact JSON format.
ΨΨΨΨΨ
    "observations_and_thoughts": "<fill in here>",
    "new_content": {
        "PLANS.txt": "<fill in here>",
        "INSIGHTS.txt": "<fill in here>"
        "price_forecasts": [
                "round": <CURRENT ROUND>,
                "min_value": 0,
                "max_value": <FORECAST UPPER BOUND>,
                "forecasted_price": "<fill in here>"
                "round": <CURRENT ROUND + 2>,
                "min_value": 0,
                "max_value": <FORECAST UPPER BOUND>,
                "forecasted_price": "<fill in here>"
                "round": <CURRENT ROUND + 5>,
                "min_value": 0,
                "max_value": <FORECAST UPPER BOUND>,
                "forecasted_price": "<fill in here>"
                "round": <CURRENT ROUND + 10>,
                "min_value": 0,
                "max_value": <FORECAST UPPER BOUND>,
                "forecasted_price": "<fill in here>"
    "submitted_orders": [
            "order_type": "<BUY or SELL>",
            "quantity": <# of STOCK units>,
            "limit_price": <LIMIT_PRICE>
            "order_type": "<BUY or SELL>",
            "quantity": <# of STOCK units>,
            "limit_price": <LIMIT_PRICE>
        // Add more or less orders as needed
B.3 Market Data Format
    Round XX:
        * Market + Portfolio Data:
            - Market price: XX
            - Market volume: XX
            - # of shares owned: XX
            - Current cash: XX
            - Stock value: XX
            - Dividend earned: XX
            - Interest earned: XX
            - Submitted orders:
                * <BUY/SELL> XX shares at XX per share
                * <BUY/SELL> XX shares at XX per share
            - Executed trades:
                -* No executed trades
        * Forecasts:
            - Round XX price forecast: XX
            - Round (XX + 2) price forecast: XX
            - Round (XX + 5) price forecast: XX
            - Round (XX + 10) price forecast: XX
B.4 Portfolio Data Format
    * Your Portfolio (Round XX):
        - Market price (Previous Round): XX
        - Buyback price: XX
        - # of shares owned: XX
        - Current cash: XX
        - Stock value: XX
Appendix C Additional Analysis
Figure 4: Human Market Trajectories
C.1 Detailed individual market analysis
  • Gemini 1.5 Pro: Gemini markets seem to generate a reverse bubble trend, where the asset is initially sold off below fundamental and then bid back up as redemption grows closer (Figure 5). While overall rationality given by the MSE ($\sim$ 3.1) from fundamental metric is low, the behavior is the exact opposite of what we have seen in the majority of human experiments and indicates that the model likely does not truly understand the task nor model any biases that humans may exhibit.
    Figure 5: Gemini 1.5 Pro Markets
    This market seems closest to the Erratic (E) hypothesis. Which is also supported by the Pearson correlation coefficient (PCC) as compared to the average human market price of -0.4009, indicating that there are large differences between this model performance and that of a humans.
    We can also analyze the uniformity of the market, as you can see in Figure 6, compared to other markets the traders do seem to have a more varied earning profile. This is also shown in the Portfolio Value variance which is: 45.58, higher than all but one of the other single-model markets.
    Figure 6: Gemini 1.5 Pro Portfolio Variance
  • Mistral Large 24.11 Mistral seems to generate markets that appear mostly like the type of bubbles we’ve seen in our human markets, though the size of the asset bubble is greatly reduced from our median human market and the price crashes below fundamental during the final rounds (Figure 7).
    Figure 7: Mistral Large 24.11 Portfolio Markets
    This behavior, while not economically rational (purely rational agents would only buy a price below 14 and would sell any price above 14) does correspond somewhat with the behavior we’ve seen out of human players, though the Pearson correlation coefficient (PCC) as compared to the average human market price of -0.1118 are large differences. This negative PCC likely pertains to the behavior in these markets after the bubble crashes. This market seems closest to the Human (H) hypothesis.
    We can also analyze the uniformity of the market, as can be seen in the below graph (Figure 8), compared to other markets the traders do seem to have a more varied earning profile. This is also shown in the mean Portfolio Value variance which is: 49.96, the highest of the single-model markets.
    Figure 8: Mistral Large 24.11 Portfolio Variance
  • Claude 3.5 Sonnet: Claude markets seem to demonstrate highly ‘rational’ behavior, the price remains very close to the fundamental value throughout the market (Figure 9). This indicates that the model likely understands the “rational” strategy. This market is the ”most” rational based on the MSE metric with the lowest MSE from fundamental of all the markets ($\sim$ 0.53).
    Figure 9: Claude 3.5 Sonnet Markets
    This market seems closest to R. Which is further supported by the Pearson correlation coefficient (PCC) as compared to the average human market price of being $\sim$ 0.0013, showing that there are large differences between this model performance and that of a humans.
    We can also analyze the uniformity of the market, as can be seen in the below graph (Figure 10), compared to other markets the traders do seem to have a more uniform earning profile. This is also shown in the mean Portfolio Value variance which is: 26.15, lower than most of the other single-model markets.
    Figure 10: Claude 3.5 Sonnet Portfolio Variance
  • GPT-3.5: Demonstrates bubble behavior without the crash, where prices rise steadily with no indication of a return to fundamental value (Figure 11). Unlike human markets, trading volume almost disappears toward the end of the experiment, whereas humans generally trade more as the end approaches. This model is the least ”rational” by the MSE metric, having a far and away the highest MSE of all the models ($\sim$ 26.36).
    Figure 11: GPT-3.5 Markets
    This market seems closest to H, as some bubbles without crashes were observed in human experiments. Interestingly, this model has a relatively high Pearson correlation coefficient (PCC) as compared to the average human market price of $\sim$ 0.4897, but it is clear that there are still large differences between this model performance and that of a humans.
    The Portfolio Value variance for this market is 30.44 (Figure 12), one of the lower values among single-model markets, indicating generally uniform behavior.
    Figure 12: GPT-3.5 Portfolio Variance
  • GPT-4o: Demonstrates mostly rational behavior, with prices remaining very close to the fundamental value throughout the market (Figure 13). This indicates that the model likely understands the “rational” strategy. This is emphasized quantitatively by the model having the second lowest MSE compared to fundamental at ($\sim$ 0.79). This market seems closest to the Rational (R) hypothesis.
    Figure 13: GPT-4o Markets
    Additionally, unlike human markets almost all the agents performed very similarly to each other, indicating that there was not a large amount of behavior variance of the agents. This is also shown in the mean Portfolio Value variance of 22.76 (Figure 14), which is the lowest amongst the single-model markets.
    Figure 14: GPT-4o Portfolio Variance
Appendix D Rationality Properties of LLM Forecasts
In addition to analyzing the pricing behavior of LLMs, we were also interested in how these models generate forecasts for future price periods. Evaluating their predictions allows us to assess not only the rationality of their forecasts—whether they align with fundamental valuation principles—but also whether the LLMs themselves behave rationally in response to their own expectations. This dual analysis helps us understand whether LLMs internalize consistent pricing logic and whether their trading behavior reflects an economically rational response to their own forecasts.
D.1 Additional Forecast Information
One challenge in adapting the experiment to LLM agents, and in doing so combining the order entry and forecasting ”screens,” was that some LLM agents struggled to adhere to the forecasting response format or input valid forecast values. We are exploring how to best mitigate this issue to ensure that our forecast data robustly captures market sentiment from all market participants. We draw a histogram to visualize the distribution of missing values in each forecast. In Figure 15, different panels represent different forecast horizons. The x-axis indicates the number of missing values per participant, while the y-axis shows the frequency. The results suggest that most agents perform well, with fewer than 10 missing forecast values. In the subsequent analysis, we exclude agents with more than 10 missing values in any forecast.
Figure 15: Histogram of Missing Forecast Valus
D.2 Forecasting Framework
Fix a session (we will suppress session notation). For each period $t$, participants provide forecasts before the price revelation. The forecast made in period $t$, denoted $f_{i,t}^{t}$, predicts the price $P_{t}$. The forecast for future periods, made in period $t$, is denoted as $f_{i,t}^{(t+h)}$, where $h\in\{0,2,5,10\}$.
For example, before prices are revealed for period $8$, participants forecast:
$f_{i,8}^{8},f_{i,8}^{10},f_{i,8}^{13},f_{i,8}^{18}$
D.3 Mean and Median Forecasts
The mean forecast for session $s$ at period $t+h$ is:
$m_{t}^{(t+h)}=\frac{\sum_{i=1}^{N_{s}}f_{i,t}^{(t+h)}}{N_{s}}$
where $N_{s}$ is the number of participants in session $s$.
The median forecast is calculated as follows: Arrange the values of $f_{t,t+h}^{(i)}$ in ascending order:
$f_{t,t+h}^{(1)},f_{t,t+h}^{(2)},\dots,f_{t,t+h}^{(N_{s})}$
where
$f_{t,t+h}^{(1)}\leq f_{t,t+h}^{(2)}\leq\dots\leq f_{t,t+h}^{(N_{s})}$
If $N_{s}$ is odd:
$\text{Median}=f_{t,t+h}^{\left(\frac{N_{s}+1}{2}\right)}$
If $N_{s}$ is even:
$\text{Median}=\frac{f_{t,t+h}^{(N_{s}/2)}+f_{t,t+h}^{(N_{s}/2+1)}}{2}$
D.4 Forecast Errors
The forecast error for individual $i$ in period $t$, forecasting $P_{t+h}$, is given by:
$E_{i,t}^{h}=P_{t+h}-f_{i,t}^{(t+h)}$
It is positive if they underforecast price and negative if they overforecast.
D.5 Summary Statistics
The Table 3 shows the Descriptive statistics and forecast errors for each model. Based on the forecast error for the current round, we find that the GPT-4o model is the closest to zero, indicating the most rational performance, with an average forecast error of 0.04. In comparison, Claude-3.5-Sonnet has an error of 0.09, Gemini-1.5-Pro at 0.14, Grok-2 at -0.17, and GPT-3.5 at -2.41.
Table 3: Descriptive statistics and forecast errors for each model.
Model Variable Mean SD Min Max Q1 Q3
claude-3.5-sonnet Price 14.65 1.00 13 20 14 15
Forecast for current round 14.67 0.98 13 20 14 15
Forecast for two rounds later 14.78 1.12 14 21 14 15
Forecast for five rounds later 14.90 1.49 14 22 14 16
Forecast for ten rounds later 15.13 2.04 14 25 14 16
Forecast error for current round 0.09 0.49 -2 3 0 0
Forecast error for two rounds later -0.01 0.68 -2 3 0 0
Forecast error for five rounds later -0.12 1.10 -4 4 -1 1
Forecast error for ten rounds later -0.38 1.52 -7 3 -1 1
gemini-1.5-pro Price 13.43 2.02 7 20 13 14.5
Forecast for current round 13.68 1.91 6 21 13 15
Forecast for two rounds later 13.87 1.77 7 22 14 15
Forecast for five rounds later 14.26 1.48 8 23 14 15
Forecast for ten rounds later 14.55 1.11 9 22 14 15
Forecast error for current round 0.14 0.88 -5 6 0 0.5
Forecast error for two rounds later -0.12 0.84 -4 4 -1 0
Forecast error for five rounds later -0.64 1.36 -5 5 -1 0
Forecast error for ten rounds later -1.23 1.69 -6 6 -2 0
gpt-3.5 Price 18.36 1.98 14 23 17.5 19.5
Forecast for current round 20.39 3.22 14 38 18 22
Forecast for two rounds later 22.35 3.31 14 42 20 24
Forecast for five rounds later 25.15 3.67 16 45 23 27
Forecast for ten rounds later 27.72 4.00 17 55 25 30
Forecast error for current round -2.41 2.44 -20 2.5 -3 -1
Forecast error for two rounds later -4.22 2.55 -24 1 -5 -3
Forecast error for five rounds later -6.76 3.15 -27 1 -8 -5
Forecast error for ten rounds later -9.05 3.78 -39 2 -11 -7
gpt-4o Price 14.94 0.87 14 20 15 15
Forecast for current round 14.95 0.91 14 21 15 15
Forecast for two rounds later 15.20 1.10 13 22 15 15
Forecast for five rounds later 15.98 1.45 14 25 15 16
Forecast for ten rounds later 17.06 2.10 14 28 16 18
Forecast error for current round 0.04 0.41 -2 2.5 0 0
Forecast error for two rounds later -0.19 0.46 -2 2 0 0
Forecast error for five rounds later -0.94 0.94 -6 1 -1 0
Forecast error for ten rounds later -2.05 1.68 -11.5 2 -3 -1
grok-2 Price 17.55 2.23 14 22 15.5 19
Forecast for current round 17.67 2.55 13 23 15 20
Forecast for two rounds later 18.18 2.94 13 25 16 20
Forecast for five rounds later 19.03 3.20 13 28 16 22
Forecast for ten rounds later 19.83 3.59 12 32 17 22
Forecast error for current round -0.17 1.62 -3 4 -1.5 1
Forecast error for two rounds later -0.43 1.48 -4 3 -2 1
Forecast error for five rounds later -0.87 1.84 -8 5 -2 0.5
Forecast error for ten rounds later -1.28 3.52 -14 7 -3 1
mistral-large Price 15.17 2.09 12 20 14 16
Forecast for current round 15.34 2.21 10 22 14 16
Forecast for two rounds later 15.77 2.45 9 23 14 17
Forecast for five rounds later 16.60 2.75 9 25 14 18
Forecast for ten rounds later 17.76 3.23 14 27 15 19
Forecast error for current round 0.00 1.04 -4 4 0 0
Forecast error for two rounds later -0.36 0.95 -4 3 -1 0
Forecast error for five rounds later -1.14 1.50 -9 4 -2 0
Forecast error for ten rounds later -2.38 2.96 -14 6 -3 0
D.6 Rationality Tests
D.6.1 Distribution of Forecast Errors
First, we examine the distribution of forecast errors of each model across different horizons, grouped by rounds. Figure 16 shows that different models have different forecast error distributions across different horizons.
At the model level, Claude-3.5-Sonnet, Gemini-1.5-Pro, and GPT-4o exhibit more centralized distributions across horizons compared to other models, with their forecasts centered around 0, indicating greater rationality. In contrast, GPT-3.5 shows the most diverse distribution. At the round level, we observe that the center of the distribution gradually shifts toward the negative for the Grok-2 and GPT-3.5 models, especially at horizons of 5 and 10, suggesting increasing bias over rounds.
Figure 16: Distribution of Forecast Errors Across Rounds
Next, we calculate the mean forecast error for each round across different horizons. Figure 17 illustrates how the mean forecast error evolves over time for each horizon. Claude-3.5-Sonnet performs the best, as its mean forecast for each horizon remains closest to 0 compared to other models. GPT-4o shows a trend of initially overforecasting prices before gradually shifting toward 0 in later rounds, indicating a movement toward greater accuracy. In contrast, GPT-3.5 performs the worst, consistently overforecasting prices as the rounds progress. Figure 18 presents a similar upward trend in the median forecast error across different horizons.
Figure 17: Mean Forecast Error by Round
Figure 18: Median Forecast Error by Round
D.6.2 Unbiasedness
At the individual agent level, we test whether the mean forecast error $E_{i,t}^{h}$ is zero across $t$ for each horizon $h$. This provides four test statistics for each value of $h$. We conduct $t$-tests to determine whether the mean forecast error is statistically different from zero.
Specifically, we test whether the mean forecast error for each model and horizon is significantly different from zero. We use the t test at agent level for four horizons and get the p value. Then we count the number of agents whose mean forecast error is significant different from 0 at each horizon, and calculate the percentage of each model. Figure 19 presents the percentage of no difference with 0 of each model at each horizon. Among the models, gpt-3.5 performs worst, as all forecast error for four horizons are all different from 0, Claude-3.5-Sonnet, Gemini-1.5-Pro performs as human, as the ratio gets lower as horizon increase. GPT-4o and Grok-2 perform best at horizon = 0, indicating rational forecasting.
Figure 19: The percentage of no difference with 0 of each model at each horizon
D.6.3 Zero Autocorrelation
To test for zero autocorrelation, we examine whether forecast errors from period $t$ should be independent of forecast errors in subsequent periods. The guiding principle is that future forecast errors should not depend on the information available at the time the forecast was made.
For example, the forecast error in period 8 is:
$E_{i,8}^{0}=P_{8}-f_{i,8}^{8}$
This error should be uncorrelated with the forecast error in period 9:
$P_{9}-f_{i,9}^{9}$
A similar zero-autocorrelation principle applies for longer horizons. To test this, we estimate the regression:
$E_{t,0}=\beta_{0}+\beta_{1}E_{t-1,0}+\epsilon_{t}$
where the hypothesis is $\beta_{1}=0$.
We divide the data of horizon $h=0$ by agent and use ordinary least squares (OLS) regression to estimate $\beta_{1}$ at the agent level for different horizons. We then count the number of agents whose $\beta_{1}$ is not significantly different from 0 ($p$-value $>0.05$) for each model and each horizon and calculate its percentage.
Figure 20 presents the results of the autocorrelation test. Among the models, GPT-3.5 appears the most rational, as it has the highest percentage of agents showing no autocorrelation. In contrast, Claude-3.5-Sonnet and Grok-2 exhibit the most autocorrelation, with the lowest percentage of agents passing the test.
Figure 20: The results of the autocorrelation test.
D.6.4 Correlation Between Forecast Errors and Forecasts
Forecast errors, $P_{t+h}-f_{i,t}^{(t+h)}$, should be uncorrelated with forecasts:
$\text{cor}(P_{t+h}-f_{i,t}^{(t+h)},f_{i,t}^{(t+h)})=0$
This principle is based on two intuitions: 1. The current forecast is part of an agent’s information set. If something in their information set can predict future errors, they should use it to reduce the error. 2. A negative correlation between forecast errors and forecasts implies that when an agent is optimistic (high $f_{i,t}^{(t+h)}$), they tend to be too optimistic (negative $P_{t+h}-f_{i,t}^{(t+h)}$).
To test this, we compute these correlations separately for each forecast horizon and agent by estimating the regression:
$E_{i,t}^{h}=\beta_{0}+\beta_{1}f_{i,t}^{(t+h)}+\epsilon_{i,t}^{h}$
for the four values of $h$. We then count the number of agents whose $\beta_{1}$ is not significantly different from zero ($p$-value $>0.05$) for each model and horizon, indicating no correlation between forecast errors and forecasts.
Figure 21 presents the results of the correlation test across different horizons and models. In this test, Gemini-1.5-Pro performs best, as it has the highest percentage across all four horizons compared to other models. In contrast, GPT-3.5 and GPT-4o perform worse, as most agents from these models exhibit forecast errors that are correlated with forecast levels.
Figure 21: Results of the correlation between forecast error and forecast